OpenStudy (anonymous):

Find the difference: (9x² − 2x + 1) − (5x² − 3)

OpenStudy (theviper):

\[(9x² − 2x + 1) − (5x² − 3)\]\[=9x² − 2x + 1-5x^2+3\]\[=4x^2-2x+4\]

OpenStudy (theviper):

Understood??

mathslover (mathslover):

ok so first solve for : \(\large{-(5x^2-3)}\)

mathslover (mathslover):

as we know that: \[\large{-(a-b)=-(a)-(-b)=-a+b}\]

mathslover (mathslover):

hence \[\large{-(5x^2-3)}\] would be equal to?

OpenStudy (anonymous):

yes understood thanks

OpenStudy (anonymous):

Change the signs of the last two terms and add the resulting five terms together.\[4 x^2-2 x+4 \]

OpenStudy (theviper):

So nice explaining @mathslover wow!!!

mathslover (mathslover):

thanks @TheViper